# Dynamical systems defined by polynomials with algebraic properties

### v4

@@ 4 Factorization of polynomials and stream zeros
We now show there are close relationships between factorization or irreducibility of a polynomial $P(z)$ and the structure of $\Omega_{P}$.
@@
Let $P(z),Q(z)$ be polynomials with integer coefficients.
(i) We have $\dim\Omega_{P}=n-h$, if $P(z)$ has the form $a_{h}z^{h}+a_{h+1}z^{h+1}+\cdots+a_{n}z^{n}$ with $a_{h}\neq 0,a_{n}\neq 0$.
(ii) If $P(z),Q(z)$ are coprime, then we have
$\{\Omega_{P}\}\cap\{\Omega_{Q}\}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$
with $\Delta=Res(P,Q)$. Therefore, $P(z),Q(z)$ are coprime if and only if $\dim(\Omega_{P}\cap\Omega_{Q})=0$.
(iii) $P(z)$ is a factor of $Q(z)$ if and only if there exists $d\in\mathbb{Z}\setminus\{0\}$ such that $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$.
(iv) If $P(z),Q(z)$ are coprime, then we have
$\Omega_{P}+\Omega_{Q}\subset\Omega_{P\times Q}\subset\frac{1}{Res(P,Q)}(\Omega_{P}+\Omega_{Q}).$
Hence, we have $\dim\Omega_{P\times Q}=\dim\Omega_{P}+\dim\Omega_{Q}$. Conversely if $P(z),Q(z)$ are not coprime, then we have $\dim\Omega_{P\times Q}>\dim\Omega_{P}+\dim\Omega_{Q}$. Specially if $Res(P,Q)=\pm 1$, then we have $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$, where $\oplus$” is the direct sum.
@@
(i) Let $k=n-h$. Since
$(P(z)\times x)_{i}=a_{h}x_{i-h}+a_{h+1}x_{i-h-1}+\cdots+a_{n}x_{i-n}$
for any $(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}$, a necessary and sufficient condition that for any $(x_{0},x_{n},\cdots,x_{l})\in\mathbb{T}^{\{0,1,\cdots,l\}}$, there exists its extension $(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ such that $P(z)\times x=(0^{\infty};0^{\infty})$ is that either $i-h$ or $i-n$ is not in $\{0,1,\cdots,l\}$ for any $i\in\mathbb{Z}$. This is because if so, then we can control the value $x_{i-h}$ or $x_{i-n}$ so that $(P(z)\times x)_{i}=0$. But if not, we have no control. This condition is equivalent to $n-h>l$. Hence, we have $\dim\Omega_{P}=n-h$.
@@
(ii) Let $P(z),Q(z)$ be coprime. Then, the resultant matrix $M$ is regular and the entries of $M^{-1}$ are in $\mathbb{Z}/\Delta$, where $\Delta=Res(P,Q)\in\mathbb{Z}\setminus\{0\}$. Since for any $k\in\mathbb{Z}$ and $\xi=(\xi_{i})_{i\in\mathbb{Z}}\in\{\Omega_{P}\}\cap\{\Omega_{Q}\}$, we have $M\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{k+n+m}\end{array}\right)\in\mathbb{Z}^{k+n+m}$, it holds that $\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{n+m}\end{array}\right)\in\frac{1}{\Delta}\mathbb{Z}^{n+m}$. This implies that $\{\Omega_{P}\}\cap\{\Omega_{Q}\}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}$, and hence, we have $\dim(\{\Omega_{P}\}\cap\{\Omega_{Q}\})=0$.
@@
Conversely, if $P(z),Q(z)$ are not coprime, then there exists a common factor of them with degree at least 1. Hence we have $\dim(\{\Omega_{P}\}\cap\{\Omega_{Q}\})\geq 1$.
@@
(iii) Assume that $P(z)$ is a factor of $Q(z)$ so that $dQ(z)=P(z)\times R(z)$, where $d\in\mathbb{Z}\setminus\{0\}$ and $R(z)$ is a polynomial with integer coefficients. Then for any $\xi\in\Omega_{P}$, we have
$dQ(z)\times\xi=(P(z)\times R(z))\times\xi=R(z)\times(P(z)\times\xi)=(0^{\infty};0^{\infty}).$
Hence, we have $\xi\in\frac{1}{d}\Omega_{Q}$ and $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$.
@@
Conversely, if $P(z)$ is not a factor of $Q(z)$, then there exists a nontrivial factor $R(z)$ of $P(z)$ such that $R(z),Q(z)$ are coprime. Then by (ii), there exists $(\xi_{i})_{i\in\mathbb{Z}}\in\Omega_{R}$ such that for any $d\in\mathbb{Z}\setminus\{0\}$, $(\xi_{i})_{i\in\mathbb{Z}}\notin\frac{1}{d}\Omega_{Q}$ holds. Since $(\xi_{i})_{\in\mathbb{Z}}\in\Omega_{P}$, $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$ does not hold for any $d\in\mathbb{Z}\setminus\{0\}$.
@@
(iv) By (iii) and the fact that $\Omega_{P\times Q}$ is a additive group, it holds that $\Omega_{P\times Q}\supset\Omega_{P}+\Omega_{Q}$. Since $P(z),Q(z)$ are coprime, there exist polynomials with integer coefficients $A(z),B(z)$ such that
$P\times A+Q\times B=\Delta,$
where $\Delta=Res(P,Q).$ Hence, we have
$P\times Q\times B=P\times(\Delta-P\times A),~{}P\times Q\times A=Q\times(\Delta-Q\times B).$
Taking $(x_{n})_{n\in\mathbb{Z}}\in\Omega_{P\times Q}$ arbitrary and by these equalities, we have
$P\times(\Delta-P\times A)\times(x_{n})_{n\in\mathbb{Z}}=Q\times(\Delta-Q\times B)\times(x_{n})_{n\in\mathbb{Z}}=(0^{\infty};0^{\infty}).$
Put
$(u_{n})_{n\in\mathbb{Z}}:=(\Delta-P\times A)\times(x_{n})_{n\in\mathbb{Z}},~{}(v_{n})_{n\in\mathbb{Z}}:=(\Delta-Q\times B)\times(x_{n})_{n\in\mathbb{Z}},$
we have $(u_{n})_{n\in\mathbb{Z}}\in\Omega_{P},~{}(v_{n})_{n\in\mathbb{Z}}\in\Omega_{Q}$ and
$(u_{n})_{n\in\mathbb{Z}}+(v_{n})_{n\in\mathbb{Z}}=(2\Delta-P\times A-Q\times B)\times(x_{n})_{n\in\mathbb{Z}}=\Delta(x_{n})_{n\in\mathbb{Z}}.$
That is, $\Delta(x_{n})_{n\in\mathbb{Z}}\in\Omega_{P}+\Omega_{Q}$, and hence, $\Omega_{P\times Q}\subset\frac{1}{\Delta}(\Omega_{P}+\Omega_{Q})$. Specially if $\Delta=\pm 1$, then we have $\Omega_{P\times Q}=\Omega_{P}+\Omega_{Q}$ and hence, by (ii), $\Omega_{P}\cap\Omega_{Q}=\{0\}$. Thus, $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$ holds.
Conversely, assume that $P(z),Q(z)$ are not coprime. Let $R(z)$ be their GCD and let $P(z)=P^{\prime}(z)R(z),~{}Q(z)=Q^{\prime}(z)R(z)$. Then, we have
$\Omega_{P^{\prime}\times R\times Q^{\prime}}\supset\Omega_{P}+\Omega_{Q}$
since $\Omega_{P^{\prime}\times R\times Q^{\prime}}$ is an additive group containing both $\Omega_{P}$ and $\Omega_{Q}$. Therefore, by (i) and the fact that $P\times Q=P^{\prime}\times R\times Q^{\prime}\times R$ we have
$\dim\Omega_{P\times Q}>\dim\Omega_{P^{\prime}\times R\times Q^{\prime}}\geq\dim(\Omega_{P}+\Omega_{Q}).$
Example 2.
Let us consider the three polynomials $Q(z)=-2z+1,~{}R(z)=-3z+1$ and $P(z)=Q(z)R(z)$. Observe that $Res(P,Q)=1$. Then, we have
$\Omega_{Q}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+1}=2x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\};$
$\Omega_{R}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+1}=3x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\};$
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+2}=5x_{n+1}-6x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\}$
and $\Omega_{Q}\cap\Omega_{R}=(0^{\infty};0^{\infty})$. Moreover, $\Omega_{P}=\Omega_{Q}\oplus\Omega_{R}$ holds.
As a consequence of Theorem 3, we have the following
@@
5 Stream zeros and the automorphism group
Let $\Omega_{P}$ be the set of stream zero’s of a primitive polynomial $P(z)$. A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is called an automorphism of $\Omega_{P}$ if it is an isomorphism of $\Omega_{P}$ as the topological additive group such that $\varphi\circ\sigma=\sigma\circ\varphi$. Furthermore, we call an automorphism $\varphi$ a strong automorphism if for any
$(x_{n};n\in\mathbb{Z}),~{}(y_{n};n\in\mathbb{Z})\in\Omega_{P}\mbox{ with }\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z}),$
$(y_{n};n\in\mathbb{Z}_{\geq 0})$ is determined by $(x_{n};n\in\mathbb{Z}_{\geq 0})$ and $(x_{n};n\in\mathbb{Z}_{\geq 0})$ is determined by $(y_{n};n\in\mathbb{Z}_{\geq 0})$. Clearly, the shift $\sigma$ is an automorphism and $I,-I$ are strong automorphisms, where ±$\pm I$ are the convolution products by ±$\pm I$.
Let $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+1\in\mathbb{Z}[z]$ with $k\geq 1,a_{k}\neq 0$. That is,
$P(z)=(0^{\infty};1,a_{1},a_{2},\cdots,a_{k-1},a_{k},0^{\infty})\in St(\mathbb{Z}).$ (3)
Therefore,
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}:~{}x_{n+k}=-a_{1}x_{n+k-1}-a_{2}x_{n+k-2}-\cdots-a_{k}x_{n}~{}(\forall n\in\mathbb{Z})\}.$
Then, we rewrite the above as
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)=M_{P}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)\mbox{ for any }n\in\mathbb{Z}\right\}.$
where $M_{P}$ denote the associate matrix as follows:

### v5

@@ 5 Factorization of polynomials and stream zeros
We now show there are close relationships between the factorization of a primitive polynomial $P(z)\in\mathbb{Z}[z]$ and a decomposition of $\Omega_{P}$ into the corresponding summands.
@@
[…]
(i) $P(z)$ and $Q(z)$ are coprime if and only if $Res(P,Q)\neq 0$. In this case, we have
$\Omega_{P}\cap\Omega_{Q}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$
with $\Delta=Res(P,Q)$. Therefore, $P(z)$ and $Q(z)$ are coprime if and only if $\Omega_{P}\cap\Omega_{Q}$ is a finite set.
(ii) $P(z)$ is a factor of $Q(z)$ if and only if $\Omega_{P}\subset\Omega_{Q}$.
(iii) If $P(z),Q(z)$ are coprime, then we have $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$, where $\oplus$” implies the semi-direct sum, that is, the intersection of the summands is a finite set.
(iv) If all the roots of $P(z)=0$ are simple, then for any $H=1,2,\cdots$, we have
$\Omega_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}.$
@@
(i) Let $P(z),Q(z)$ be coprime. Then, the resultant matrix $\mathcal{R}$ is regular and the entries of $\mathcal{R}^{-1}$ are in $\mathbb{Z}/\Delta$, where $\Delta=Res(P,Q)\in\mathbb{Z}\setminus\{0\}$. Since for any $k\in\mathbb{Z}$ and $\xi=(\xi_{i})_{i\in\mathbb{Z}}\in\Omega_{P}\cap\Omega_{Q}$, we have $\mathcal{R}\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{k+n+m}\end{array}\right)\in\mathbb{Z}^{k+n+m}$, it holds that $\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{n+m}\end{array}\right)\in\frac{1}{\Delta}\mathbb{Z}^{n+m}$. This implies that $\Omega_{P}\cap\Omega_{Q}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$, and hence, $\Omega_{P}\cap\Omega_{Q}$ ia a finite set by Lemma 3.
@@
Conversely, if $P(z),Q(z)$ are not coprime, then there exists a common factor of them with degree at least 1. Hence we have $\dim(\Omega_{P}\cap\Omega_{Q})\geq 1$, and hence, $\Omega_{P}\cap\Omega_{Q}$ is an infinite set.
@@
(ii) Assume that $P(z)$ is a factor of $Q(z)$ so that $Q(z)=P(z)\times R(z)$, where $R(z)$ is a polynomial with integer coefficients. Then for any $\xi\in\Omega_{P}$, we have
$Q(z)\times\xi=(P(z)\times R(z))\times\xi=R(z)\times(P(z)\times\xi)=(0^{\infty};0^{\infty}).$
Hence, we have $\xi\in\Omega_{Q}$ and $\Omega_{P}\subset\Omega_{Q}$.
@@
Conversely, if $P(z)$ is not a factor of $Q(z)$, then there exists a nontrivial factor $R(z)$ of $P(z)$ such that $R(z),Q(z)$ are coprime. Then by (i), $\Omega_{Q}\cap\Omega_{R}$ is a finite set, while $\Omega_{R}$ is not. Hence, $\Omega_{R}\subset\Omega_{Q}$ does not hold. Thus, $\Omega_{P}\subset\Omega_{Q}$ does not hold since $\Omega_{P}\supset\Omega_{R}$.
(iii) By (ii) and the fact that $\Omega_{P\times Q}$ is a additive group, it holds that
$\Omega_{P\times Q}\supset\Omega_{P}+\Omega_{Q}.$ (7)
@@
Since $P(z),Q(z)$ are coprime, there exist polynomials with integer coefficients $A(z),B(z)$ with $\deg A(z)\leq h$ and $\deg B(z)\leq k$ such that
$P\times A+Q\times B=\Delta,$ (8)
where $\Delta=Res(P,Q)$. By (8), we have
$P\times Q\times B=P\times(\Delta-P\times A),~P\times Q\times A=Q\times(\Delta-Q\times B).$ (9)
Take $X:=(x_{n};n\in\mathbb{Z})\in\Omega_{P\times Q}$ arbitrary. By Corollary 1, there exists $Y:=(y_{n};n\in\mathbb{Z})\in\Omega_{P\times Q}$ such that $\Delta Y=X$. Since $Y\in\Omega_{P\times Q}$, we have by (9) that
$P\times(\Delta-P\times A)\times Y=Q\times(\Delta-Q\times B)\times Y=(0^{\infty};0^{\infty}).$
Let
$(u_{n};n\in\mathbb{Z})=(\Delta-P\times A)\times Y,~(v_{n};n\in\mathbb{Z})=(\Delta-Q\times B)\times Y.$
Then, from above $(u_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $(v_{n};n\in\mathbb{Z})\in\Omega_{Q}$. Moreover, by (8),
$(u_{n};n\in\mathbb{Z})+(v_{n};n\in\mathbb{Z})=(2\Delta-P\times A-Q\times B)\times Y=\Delta Y=X.$
Thus, we have $\Omega_{P\times Q}\subset\Omega_{P}+\Omega_{Q}$, hence together with (7) and (i), we complete the proof.
(iv) Considering the lifts, the equation
$\Omega_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}$
follows from the equation
$\tilde{\Omega}_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\tilde{\Omega}_{P}.$
For any $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$, define the derivative $X^{\prime}\in St(\mathbb{R})$ to be $\sum_{n\in\mathbb{Z}}nx_{n}z^{n-1}$ as usual. That is, $X^{\prime}=\sigma(\mathbf{n}\odot X)$, where $\sigma$ is the shift on $St(\mathbb{T})$. Clearly, if $X\in\Omega_{P}$, then $(P\times X)^{\prime}=(0^{\infty};0^{\infty})$.
Hence if $X\in\Omega_{P^{H-1}}$, then we have
$(0^{\infty};0^{\infty})=(P^{H}\times X)^{\prime}=HP^{H-1}P^{\prime}\times X+P^{H}\times X^{\prime}=P^{H}\times\sigma(\mathbf{n}\odot X).$
Therefore, $\sigma(\mathbf{n}\odot X)\in\Omega_{P^{H}}$ and $\mathbf{n}\odot X\in\Omega_{P^{H}}$. Thus, $\mathbf{n}\odot\Omega_{P^{H-1}}\subset\Omega_{P^{H}}$. It follows from this that
$\Omega_{P^{H}}\supset\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}~~(H=1,2,\cdots).$ (10)
To prove the oposite inclusion, we use the local lifts of these terms. By Corollary 1,
$\tilde{\Omega}_{P^{H}}|_{N}\supset\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\tilde{\Omega}_{P}|_{N}$
holds for any large $N\in\mathbb{Z}_{>0}$. Since the both sides are vector spaces over $\mathbb{R}$, they coincide if both sides have the same dimension $kH$ as that of the vector spaces over $\mathbb{R}$.
For this purpose, we need the following lemma which is a generalization of the Vandermonde’s determinant due to R. P. Flowe and A. G. Harris [4].
Lemma 4.
[4] Let $l_{1},\cdots,l_{h}$ and $L$ be positive integers such that $l_{1}+\cdots+l_{h}=L$. Let
$\mathcal{M}[x_{1},\cdots,x_{h};L;l_{1},\cdots,l_{h}]:=\left(\mathcal{M}[x_{1},L,l_{1}]\mathcal{M}[x_{2},L,l_{2}]\cdots\mathcal{M}[x_{h},L,l_{h}]\right)$
be the $L\times L$-matrix whose columns from 1st to $l_{1}$-th is $\mathcal{M}[x_{1},L,l_{1}]$ and from $l_{1}+1$-th to $l_{1}+l_{2}$-th is $\mathcal{M}[x_{2},L,l_{2}]$, $\cdots$, where $x_{1},x_{2},\cdots,x_{h}\in\mathbb{C}$ and
$\mathcal{M}(x,L,l)=\left(\begin{array}[]{ccccc}1&0&0&\cdots&0\\ x&x&x&\cdots&x\\ x^{2}&2x^{2}&2^{2}x^{2}&\cdots&2^{l-1}x^{2}\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ x^{L-1}&(L-1)x^{L-1}&(L-1)^{2}x^{L-1}&\cdots&(L-1)^{l-1}x^{L-1}\end{array}\right).$
Then
$\det\mathcal{M}[x_{1},\cdots,x_{h};L;l_{1},\cdots,l_{h}]=C\prod_{i=1}^{h}x_{i}^{l_{i}(l_{i}-1)/2}\prod_{i<j}(x_{j}-x_{i})^{l_{i}l_{j}}$ (11)
with $C=\prod_{i=1}^{h}\prod_{j=1}^{l_{i}}j!\neq 0$.
Continuation of the proof of (iv) of Theorem 4:  Let $x_{1},x_{2},\cdots,x_{s}$ be distinct real numbers and
$y_{1},y_{2},\cdots,y_{u};\overline{y}_{1},\overline{y}_{2},\cdots,\overline{y}_{u}$
be distinct complex numbers not being real which are the set of roots of $P(z)=0$, where $\overline{y}$ is the complex conjugate of $y$. Then, $s+2u=k$ and the roots of $P(z)^{H}=0$ are as above with the multiplicity $H$. Hence, by Lemma 2, the set of vectors
$\{\mathbf{n}^{\odot i}(x_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,s\}$
$\cup$ $\{\mathbf{n}^{\odot i}(y_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,u\}$
$\cup$ $\{\mathbf{n}^{\odot i}(\overline{y}_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,u\}$
spans $kH$-dimensional vector space over $\mathbb{C}$. Replacing the pair
$(y_{i}^{n};n\in\mathbb{Z}),~(\overline{y}_{i}^{n};n\in\mathbb{Z})~\mbox{by}~(\Re(y_{i}^{n});n\in\mathbb{Z}),~(\Im(y_{i}^{n});n\in\mathbb{Z}),$
we get the same space whose restriction to the real parts is spanned by the replaced one. This vector space over $\mathbb{R}$ is same as the vector space spanned by
$\tilde{\Omega}_{P},~\mathbf{n}\odot\tilde{\Omega}_{P},~\cdots,~\mathbf{n}^{\odot(H-1)}\tilde{\Omega}_{P},$
and hence, has the dimension $kH$ which coincides the dimension of the space $\tilde{\Omega}_{P^{H}}$. Thus, we proved the oposite inclusion of (10). ∎
The following Theorem is already known in S. Akiyama, T. Kamae & H. Kaneko [8] except for the statement of semi-direct.
Theorem 5.
Let $P(z)\in\mathbb{Z}[z]$ be a primitive polynomial with the resolution into irreducible factors
$P(z)=Q_{1}(z)^{r_{1}}\cdots Q_{h}(z)^{r_{h}},$
where $Q_{i}(z)\in\mathbb{Z}[z]~(i,1,\cdots,h)$. Then, we have
$\Omega_{P}=\sum_{i=1}^{k}\sum_{j=1}^{r_{i}-1}\mathbf{n}^{\odot j}\odot\Omega_{Q_{j}},$
where the sum is semi-direct.
Proof.
By (iii) and (iv) of Theorem 4, we have
$\Omega_{P}=\sum_{i=1}^{k}\Omega_{Q_{i}}^{r_{i}}$
$=\sum_{i=1}^{k}\sum_{j=1}^{r_{i}-1}\mathbf{n}^{\odot j}\odot\Omega_{Q_{j}}.$
As a consequence of Theorem 5, we have the following Corollary.
@@
6 Stream zeros and the strong automorphism group
Let
$P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+a_{0}\in\mathbb{Z}[z]~(k\geq 1,a_{0}\neq 0,a_{k}\neq 0)$ (12)
be a primitive polynomial and